# SLIM: Scalable and Lightweight LiDAR Mapping in Urban Environments

### v1

@@ Abstract
LiDAR point cloud maps are extensively utilized on roads for robot navigation due to their high consistency. However, dense point clouds face challenges of high memory consumption and reduced maintainability for long-term operations. In this study, we introduce SLIM, a scalable and lightweight mapping system for long-term LiDAR mapping in urban environments. The system begins by parameterizing structural point clouds into lines and planes. These lightweight and structural representations meet the requirements of map merging, pose graph optimization, and bundle adjustment, ensuring incremental management and local consistency. For long-term operations, a map-centric nonlinear factor recovery method is designed to sparsify poses while preserving mapping accuracy. We validate the SLIM system with multi-session real-world LiDAR data from classical LiDAR mapping datasets, including KITTI, NCLT, and HeLiPR. The experiments demonstrate its capabilities in mapping accuracy, lightweightness, and scalability. Map re-use is also verified through map-based robot localization. Ultimately, with multi-session LiDAR data, the SLIM system provides a globally consistent map with low memory consumption (130 KB/km). We have made our code open-source to benefit the community.
@@
A video is submitted as a multimedia attachment.
@@ I Introduction
The urban environment is the primary setting for our daily lives. Mapping in urban environments has a rich history, dating back to the development of autonomous vehicles during the DARPA Urban Challenge [3]. With advancements in LiDAR SLAM, mobile robots are now well-developed to handle mapping tasks in well-structured urban areas. However, deploying existing systems on mobile robots for long-term use presents several challenges in urban environments, stated as follows
@@
    We introduce a map-centric nonlinear graph sparsification method to manage map size as sessions increase, ensuring scalability for long-term maintenance.
@@
    We validate the SLIM system on three different multi-session datasets in terms of accuracy, lightweightness, and scalability.
The proposed SLIM system is an extended version of our previous work in [8]. This enhanced version introduces improvements across three key aspects. Firstly, it no longer relies on semantic information for line and plane extraction, focusing solely on geometric characteristics. We believe that additional semantic information from visual perception can enhance SLIM’s performance in practice. Secondly, the map merge module has been redesigned to improve robustness and efficiency for urban mapping. Lastly, we introduce a map-centric nonlinear factor recovery method to ensure the system’s scalability. In summary, these improvements make the SLIM system more applicable for long-term urban mapping.
Figure 1: Demonstration of SLIM on two regions in the HeLiPR dataset [9]. Best viewed zoomed in and in color. SLIM provides parameterized maps with lines (colored in cyan) and planes (colored in magenta). We also display downsampled point cloud maps for comparison (in light blue). The SLIM maps are naturally more lightweight than conventional LiDAR point cloud maps. The map consumption is significantly reduced from 239.50 MB to 1.20 MB in (a), and from 82.38 MB to 0.72 MB in (b).
The rest of our paper is organized as follows: Section II presents related work on LiDAR mapping approaches. Section III provides an overview of the designed SLIM system. Section IV details the representations, map merge, and optimization within the SLIM system. To address scalability, Section V introduces our proposed map-centric nonlinear factor recovery for long-term operation. We validate the proposed system in real-world scenarios in Section VI. Discussions and limitations are presented in Section VII for in-depth analysis. The findings and future studies are summarized in Section VIII.
@@ II Related Work
LiDAR mapping techniques have a long history, dating back twenty years, with LiDAR SLAM being the most widely used approach. This section first focuses on multi-session LiDAR SLAM, a prominent research topic over the past five years. Subsequently, we discuss the representations of LiDAR SLAM and long-term mapping techniques.
@@ II-A Multi-session LiDAR SLAM
Research on LiDAR SLAM has achieved significant success. Representative works include LOAM [10], Fast-LIO2 [4], DLO [11], and BALM2 [12]. These can be categorized into two types: the majority [4, 11, 10, 13, 14, 15] adopt a scan-to-map scheme with a fixed map, demonstrating high efficiency and accuracy. The second branch is BA-based methods, which parameterize map points as planes [16, 12, 17, 18], occupancy fields [19], and implicit maps [20]. BA-based methods optimize the map along with the pose to achieve higher accuracy compared to scan-to-map approaches. As LiDAR SLAM has developed, researchers have shifted focus towards multi-session LiDAR SLAM. This involves assembling individual maps and poses from multiple sessions to construct a more complete and accurate globally consistent map, under either a distributed or centralized scheme.
@@
The general pipeline for assembling multi-session maps in different coordinates involves two key steps: loop closure detection and false loop pruning [21]. These steps aim to establish correspondences and eliminate incorrect ones, respectively. Several methods employ different approaches for loop closure detection. Maplab [22, 23] and DiSCo-SLAM [24] utilize the Scan Context descriptor [25] to detect loops between maps. However, handcrafted methods may struggle with significant viewpoint changes and have limited descriptor capabilities. In contrast, AutoMerge [26] uses a more powerful deep learning-based method, an improved version of FusionVLAD [27], to detect loops in large-scale environments. To prune false loops, various heuristics are introduced to reject potential false positives, including quality assessment of point cloud registration [28], loop closure prioritization [29], and ensuring sequential consistency [30]. While heuristic methods provide a rapid means of rejecting obvious false loops, they can still suffer from appearance ambiguities. Therefore, most multi-session LiDAR SLAM approaches [24, 31, 32, 33] employ pairwise consistent measurement (PCM) set maximization [34] to prune outliers. This involves finding the maximum geometrically consistent loop set, thereby decreasing false loop associations with similar appearances.
@@
To ensure the global consistency of the merged map, PGO is utilized to align trajectories from individual sessions and mitigate accumulated drift from LiDAR odometry. Latif et al. [35] introduce a novel algorithm, Realizing, Reversing, and Recovering (RRR), to identify and reject loops that violate odometry constraints. To improve efficiency, LTA-OM [36] eliminates the need for re-estimating graph parameters by evaluating the residual outcomes of key point factors and determining whether to recover the backup graph based on these residuals. However, these methods, which involve the incremental addition and rejection of loops, face efficiency limitations due to the requirement of multiple optimization rounds. To address this issue, contemporary multi-session SLAM systems, such as Kimera-Multi [32] and LAMP 2.0 [37], incorporate all loop factors into the M-Estimator optimization to reject outliers. Subsequently, Graduated Non-Convexity (GNC) [38] is employed to achieve robust PGO. In contrast to the aforementioned approaches that focus on global consistency using PGO, HBA [39] employs LiDAR BA to enhance map consistency directly and introduces a hierarchical structure to reduce computation cost.
@@ II-B2 Implicit representation
Implicit representations in radiance [52], distance [53], and occupancy [54] fields have significantly influenced LiDAR SLAM research. SHINE-Mapping [55] presents an incremental implicit mapping system incorporating online learning of sparse octree node features. LONER [56] adopts a hierarchical feature grid encoding inspired by Ins-NGP, building a real-time neural implicit LiDAR SLAM system enhanced with LiDAR odometry. NeRF-LOAM [57] introduces a dynamic voxel embedding generation approach designed for large-scale scenario. NeRF-LOAM incorporates a neural front end to minimize SDF errors from a fixed implicit network. These implicit neural LiDAR SLAM methods exhibit impressive mapping accuracy. However, they either necessitate mesh map reconstruction for localization or depend directly on the implicit map, constraining their operational frequency for real-time use and long-term scalable mapping.
@@ II-C Long-term Mapping
Long-term mapping presents challenges for visual-based frameworks due to illumination changes [58, 59]. Unlike visual sensing, LiDAR sensing remains consistent and illumination-invariant across seasons, making it well-suited for long-term mapping. At the front end, managing newly generated maps from robots has been extensively studied. Typically, map merging involves place recognition and point cloud registration techniques [21]. For long-term managment, compressing point cloud size can reduce computational demands for online applications, utilizing observation-based [60] and geometric-based sampling [61].
@@ III System Overview
The system pipeline is illustrated in Figure 2. Given sequential input LiDAR scans, SLIM first converts raw LiDAR point clouds into line and plane representations based on geometric properties. This vectorization pipeline is detailed in Section IV-A. These parameterized representations are compact and manageable for subsequent tasks. Lines and planes are accumulated using LiDAR odometry or with additional onboard odometry, resulting in lightweight landmarks in maps. To merge multi-session maps, SLIM utilizes place recognition followed by local pose estimation, with the map merge detailed in Section IV-B. Though the maps are unified in the same coordinate, the drift from odometry still exists. In Section IV-C, SLIM achieves map refinement (smoothing) through a coarse-to-fine approach: PGO and BA, designed based on the parameterized lines and planes. With increased merged map sessions, a map-centric nonlinear factor recovery is implemented to keep the number of poses manageable, as presented in Section V. All these operations mentioned above make the mapping pipeline scalable with increasing multi-session LiDAR maps. We also summarize the pipeline in Algorithm 1.
It is worth noting that the map landmarks consist solely of parameterized lines and planes after map vectorization, without storing any dense point clouds in the map. Thus, the back-end map optimization and maintenance utilize the same representations. In the experimental section, we validate that these lightweight representations effectively support long-term urban mapping deployment with various mobile robots equipped with different 3D LiDAR sensors.
@@
Algorithm 1 Pipeline of SLIM
@@ IV Map Representation, Merge and Refinement
In this section, we first present the front-end vectorization of the SLIM system, which converts raw LiDAR point clouds into parameterized lines and planes. Subsequently, the map merge function is introduced to unify local maps within the global frame through place recognition and global registration. Map refinements enhance mapping accuracy using our proposed PGO and BA techniques.
@@ IV-A1 Pre-processing and Feature Points Selection
Given sequential LiDAR scans from robots, we apply existing LiDAR odometry methods, such as LOAM [10] and KISS-ICP [14], to obtain local maps and sequential robot poses. Sensor fusion approaches [4] could also be an alternative. These methods provide dense point clouds and robot poses with drifts, serving as input for the SLIM system. Typically, robots or vehicles operate in different parts of urban environments, corresponding to different local frames. The SLIM system is desired the input data into a single global frame and generates consistent and lightweight mapping results with minimal drift.
@@
The most common objects in urban environments are buildings, roads, and roadside poles, which can be categorized into two types of features: lines and planes. For line feature extraction, we initially divide points into their respective scan lines to restore the original scanning pattern. It is important that points on the same scan line are arranged in the order of scanning time. Subsequently, we calculate the bilateral distance difference for a point on a scan line. Based on the distance difference values on both sides, we label points as negative gradient points, positive gradient points, and points with gradients on both sides using a threshold. We then identify pairs of neighboring negative and positive gradient points that are sufficiently close in distance. Extracting the points between them, as well as those with gradients on both sides, allows us to obtain a candidate point set. Finally, we employ clustering algorithms and RANSAC line fitting methods to refine this candidate point set, ultimately obtaining multiple line segments. Planar points are selected based on the local structure of points. First, TRAVEL [79] is used to segment the ground from point clouds, identifying ground and non-ground points. Hash octrees are then constructed separately for these two different types of points. Classical principal component analysis (PCA) is applied to identify planar points, treating them as plane segments. These line and plane segments are significantly less than the raw map points, preserving the original geometric information and making them advantageous for map management in large-scale urban environments.
@@
In addition to segment extractions, we introduce information matrices to describe uncertainties. We define $\sqrt{\mathbf{\Lambda}}=\mathbf{I}\cdot(\sqrt{N}/\sigma)$ as the square root of the information matrix, where $(\sqrt{\mathbf{\Lambda}})^{T}\sqrt{\mathbf{\Lambda}}=\mathbf{\Lambda}$. The term $N$ represents the number of points in line or plane segments, and the variance $\sigma$ varies across different semantic categories. We set $\sigma_{r}=0.1,\sigma_{b}=0.2,\sigma_{p}=0.3$ for roads, buildings, and poles, respectively. The line observation is defined as $\mathbf{f}_{\mathcal{L}}:=\langle\mathbf{\hat{p}}_{a},\mathbf{\hat{p}}_{b},\sqrt{\mathbf{\Lambda}_{\mathcal{L}}},N\rangle$, and the plane observation is defined as $\mathbf{f}_{\mathcal{S}}:=\langle\mathbf{\hat{p}}_{a},\mathbf{\hat{p}}_{b},\mathbf{\hat{p}}_{c},\sqrt{\mathbf{\Lambda}_{\mathcal{S}}},N\rangle$. Two observation points serve as the residuals for line landmarks, while three are used for plane landmarks. To ensure that the selected sparse observation points represent the local point cloud structure, we base their selection on the eigenvalue decomposition results of local PCA.

### v2

@@ Abstract
LiDAR point cloud maps are extensively utilized on roads for robot navigation due to their high consistency. However, dense point clouds face challenges of high memory consumption and reduced maintainability for long-term operations. In this study, we introduce SLIM, a scalable and lightweight mapping system for long-term LiDAR mapping in urban environments. The system begins by parameterizing structural point clouds into lines and planes. These lightweight and structural representations meet the requirements of map merging, pose graph optimization, and bundle adjustment, ensuring incremental management and local consistency. For long-term operations, a map-centric nonlinear factor recovery method is designed to sparsify poses while preserving mapping accuracy. We validate the SLIM system with multi-session real-world LiDAR data from classical LiDAR mapping datasets, including KITTI, NCLT, HeLiPR and M2DGR. The experiments demonstrate its capabilities in mapping accuracy, lightweightness, and scalability. Map re-use is also verified through map-based robot localization. Finally, with multi-session LiDAR data, the SLIM system provides a globally consistent map with low memory consumption ($\sim$130 KB/km on KITTI).
@@
A video is submitted as a multimedia attachment. A public version can be accessed online. Code is released at https://github.com/HKUST-Aerial-Robotics/SLIM.
@@ I Introduction
The urban environment is the primary setting of our daily lives. Mapping in urban environments has a rich history, dating back to the development of autonomous vehicles during the DARPA Urban Challenge [3]. With advancements in LiDAR SLAM, mobile robots are now well-developed to handle mapping tasks in well-structured urban areas. However, deploying existing systems on mobile robots for long-term use presents several challenges in urban environments, stated as follows
@@
    We introduce a map-centric marginalization scheme to manage map size as sessions increase, ensuring scalability for long-term maintenance. The key contribution is an efficient nonlinear factor recovery (NFR) method based on parameterized landmarks.
@@
    We validate the SLIM system on four different multi-session datasets in terms of accuracy, lightweightness, and scalability.
The proposed SLIM system is an extended version of our previous work in [8]. This enhanced version introduces improvements across three key aspects. Firstly, it no longer relies on semantic information for line and plane extraction, focusing solely on geometric characteristics. We believe that additional semantic information from visual perception can enhance the mapping performance in practice. Secondly, the map merge module has been redesigned to improve robustness and efficiency for urban mapping. Lastly, we introduce a map-centric nonlinear factor recovery method to improve the system’s scalability. In summary, these improvements make the SLIM system more applicable for long-term urban mapping.
Figure 1: Demonstration of SLIM using the HeLiPR dataset [9]. Best viewed zoomed in and in color. SLIM provides parameterized maps with lines (colored in cyan) and planes (colored in magenta). For comparison, we also present point clouds (in light blue), which are widely used in conventional LiDAR mapping systems. More mapping results are presented in Figure 10. Furthermore, we compare the memory consumption of SLIM-generated maps with point cloud maps of varying densities (downsampled with radius $r$). In this HeLiPR session, SLIM maps are inherently more lightweight than LiDAR point cloud maps, with memory usage significantly reduced from 690.60 MB (point clouds) to just 0.41 MB (SLIM).
The rest of our paper is organized as follows: Section II presents related work on LiDAR mapping approaches. Section III provides an overview of the designed SLIM system. Section IV details the representations, map merge, and optimization within the SLIM system. Section V introduces our proposed map-centric marginalization for long-term operation. We validate the proposed system in real-world scenarios in Section VI. Discussions and limitations are presented in Section VII for in-depth analysis. The findings and future studies are summarized in Section VIII.
@@ II Related Work
This section first focuses on multi-session LiDAR SLAM, a prominent research topic over the past five years. Subsequently, we discuss the representations of LiDAR SLAM and long-term mapping techniques.
@@ II-A Multi-session LiDAR SLAM
Research on LiDAR SLAM has achieved significant success. Representative works include LOAM [10], Fast-LIO2 [4], DLO [11], and BALM2 [12]. These can be categorized into two types: the majority [4, 11, 10, 13, 14] adopt a scan-to-map scheme with a fixed map, demonstrating high efficiency and accuracy. The second branch is BA-based methods, which parameterize map points as planes [15, 12, 16, 17], occupancy fields [18], and implicit maps [19]. BA-based methods optimize the map along with the pose to achieve higher accuracy compared to scan-to-map approaches. As LiDAR SLAM has developed, researchers have shifted focus towards multi-session LiDAR SLAM. This involves assembling individual maps and poses from multiple sessions to construct a more complete and accurate globally consistent map, under either a distributed or centralized scheme.
@@
The general pipeline for assembling multi-session maps in different coordinates involves two key steps: loop closure detection and false loop pruning [20]. These steps aim to establish correspondences and eliminate incorrect ones, respectively. Several methods employ different approaches for loop closure detection. Maplab [21] and DiSCo-SLAM [22] utilize the Scan Context descriptor [23] to detect loops between maps. However, handcrafted methods may struggle with significant viewpoint changes and have limited descriptor capabilities. In contrast, AutoMerge [24] uses a more powerful deep learning-based method, an improved version of FusionVLAD [25], to detect loops in large-scale environments. To prune false loops, various heuristics are introduced to reject potential false positives, including quality assessment of point cloud registration [26], loop closure prioritization [27], and ensuring sequential consistency [28]. While heuristic methods provide a rapid means of rejecting obvious false loops, they can still suffer from appearance ambiguities. Therefore, most multi-session LiDAR SLAM approaches [22, 29, 30, 31] employ pairwise consistent measurement (PCM) set maximization [32] to prune outliers. This involves finding the maximum geometrically consistent loop set, thereby decreasing false loop associations with similar appearances.
@@
To ensure the global consistency of the merged map, PGO is employed to align the trajectories from individual sessions and to reduce the accumulated drift caused by LiDAR odometry. Latif et al. [33] introduce a novel algorithm, Realizing, Reversing, and Recovering (RRR), to identify and reject loops that violate odometry constraints. To improve efficiency, LTA-OM [34] eliminates the need for re-estimating graph parameters by evaluating the residual outcomes of key point factors and determining whether to recover the backup graph based on these residuals. However, these methods, which involve the incremental addition and rejection of loops, face efficiency limitations due to the requirement of multiple optimization rounds. To address this issue, contemporary multi-session SLAM systems, such as Kimera-Multi [30] and LAMP 2.0 [35], incorporate all loop factors into the M-Estimator optimization to reject outliers. Subsequently, Graduated Non-Convexity (GNC) [36] is employed to achieve robust PGO. In contrast to the aforementioned approaches that focus on global consistency using PGO, HBA [37] employs LiDAR BA to enhance map consistency directly and introduces a hierarchical structure to reduce computation cost.
@@ II-B2 Implicit representation
Implicit representations in radiance [50], distance [51], and occupancy [52] fields have significantly influenced LiDAR SLAM research. SHINE-Mapping [53] presents an incremental implicit mapping system incorporating online learning of sparse octree node features. LONER [54] adopts a hierarchical feature grid encoding inspired by Ins-NGP, building a real-time neural implicit LiDAR SLAM system enhanced with LiDAR odometry. NeRF-LOAM [55] introduces a dynamic voxel embedding generation approach designed for large-scale scenario. These implicit neural LiDAR SLAM methods exhibit impressive mapping accuracy. However, they either necessitate mesh map reconstruction for localization or depend directly on the implicit map, constraining their operational frequency for real-time use and long-term scalable mapping.
@@ II-C Long-term Mapping
Long-term mapping presents challenges for visual-based frameworks due to illumination changes [56, 57]. Unlike visual sensing, LiDAR sensing remains consistent and illumination-invariant across day-night variations, making it well-suited for long-term mapping. At the front end, managing newly generated maps from robots has been extensively studied. Typically, map merging involves place recognition and point cloud registration techniques [20]. For long-term managment, compressing point cloud size can reduce computational demands for online applications, utilizing observation-based [58] and geometric-based sampling [59].
@@ III System Overview
The system pipeline is illustrated in Figure 2. Given sequential input LiDAR scans, SLIM first converts raw LiDAR point clouds into line and plane representations based on geometric properties. This vectorization pipeline is detailed in Section IV-A. These parameterized representations are compact and manageable for subsequent tasks. Lines and planes are accumulated using LiDAR odometry or with additional onboard odometry, resulting in lightweight landmarks in maps. To merge multi-session maps, SLIM utilizes a global registration approach without place recognition, detailed in Section IV-B. Though the maps are unified in the same coordinate, the drift from odometry still exists. In Section IV-C, SLIM achieves map refinement (smoothing) through a coarse-to-fine approach: PGO and BA, designed based on the parameterized lines and planes. With increased merged map sessions, a map-centric nonlinear factor recovery is implemented to keep the number of poses manageable, as presented in Section V. All these operations mentioned above make the mapping pipeline scalable with increasing multi-session LiDAR maps. We also summarize the pipeline in Algorithm 1. It is worth noting that the map landmarks consist solely of parameterized lines and planes after map vectorization, without storing any dense point clouds in the map.
@@
Algorithm 1 SLIM Pipeline
@@ IV Map Representation, Merge and Refinement
We first present the front-end map vectorization in Section IV-A. Subsequently, the map merge function is introduced in Section IV-B to unify local maps within the global frame through place recognition and global registration. In Section IV-C, map refinement modules improve mapping accuracy using our proposed PGO and BA approaches.
@@ IV-A1 Pre-processing and Feature Points Selection
Given sequential LiDAR scans from robots, we apply existing LiDAR odometry methods, such as LOAM [10] and KISS-ICP [14], to obtain local maps and sequential robot poses. Sensor fusion approaches [4] could also be an alternative. These methods provide dense point clouds and robot poses with drifts, serving as input for the SLIM system. Typically, robots or vehicles operate in different parts of urban environments, corresponding to different local frames. The SLIM system is designed to provide a consistent and lightweight map in a global frame.
@@
The most common objects in urban environments are buildings, roads, and roadside poles, which can be categorized into two types of features: lines and planes. For line feature extraction, we initially divide points into their respective scan lines to restore the original scanning pattern. It is important that points on the same scan line are arranged in the order of scanned timestamp. Subsequently, we calculate the bilateral distance difference for a point on a scan line. Based on the distance difference values on both sides, we label points as negative gradient points, positive gradient points, and points with gradients on both sides with a threshold. We then identify pairs of neighboring negative and positive gradient points that are sufficiently close in distance. Extracting the points between them, as well as those with gradients on both sides, allows us to obtain a candidate point set. Finally, we employ clustering and line fitting methods to refine this candidate point set, finally obtaining multiple line segments. Planar points are selected based on the local structure of points. First, TRAVEL [77] is used to identify ground and non-ground points. Hash octrees are then constructed separately for these two different types of points. Classical principal component analysis (PCA) is applied to identify planar points and convert them to plane segments. These line and plane segments are significantly fewer than the raw LiDAR points while preserving the original geometric information.
We define a line observation as $\mathbf{f}_{\mathcal{L}}:=\langle\mathbf{\hat{p}}_{a},\mathbf{\hat{p}}_{b},\sqrt{\mathbf{\Lambda}_{\mathcal{L}}},N\rangle$, and a plane observation as $\mathbf{f}_{\mathcal{S}}:=\langle\mathbf{\hat{p}}_{a},\mathbf{\hat{p}}_{b},\mathbf{\hat{p}}_{c},\sqrt{\mathbf{\Lambda}_{\mathcal{S}}},N\rangle$. Two observation points are used to build residuals for line landmarks, while three are used for plane landmarks. To ensure that these selected sparse observation points effectively represent the local point cloud structure, their selection is based on the eigenvalue decomposition results of a local PCA. More specifically, we assume that all extracted LiDAR points of a single observation is $\left\{\mathbf{p}_{i}\right\}$ and their centroid is $\mathbf{\bar{p}}$. The computed eigenvalues are $\lambda_{1}<\lambda_{2}<\lambda_{3}$ and corresponding eigenvectors are $\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3}$. The two line observation points are defined as
$\mathbf{\hat{p}}_{a}$ $=\mathbf{\bar{p}}+\sqrt{2\lambda_{3}}\mathbf{v}_{3}$ (1)
$\mathbf{\hat{p}}_{b}$ $=\mathbf{\bar{p}}-\sqrt{2\lambda_{3}}\mathbf{v}_{3}$
Similarly, the three plane observation points are defined as
$\mathbf{\hat{p}}_{a}$ $=\mathbf{\bar{p}}+\sqrt{2\lambda_{2}}\mathbf{v}_{2}$ (2)
$\mathbf{\hat{p}}_{b}$ $=\mathbf{\bar{p}}-\sqrt{\frac{\lambda_{2}}{2}}\mathbf{v}_{2}+\sqrt{\frac{\lambda_{1}}{2}}\mathbf{v}_{1}$
$\mathbf{\hat{p}}_{c}$ $=\mathbf{\bar{p}}-\sqrt{\frac{\lambda_{2}}{2}}\mathbf{v}_{2}-\sqrt{\frac{\lambda_{1}}{2}}\mathbf{v}_{1}$
It is worth noting that the degrees-of-freedom (DoF) for the line residuals of two observation points are 4, which is consistent with the DoF of line landmarks. Similarly, the DoF for the plane residuals of three observation points are 3, which is consistent with the DoF of plane landmarks. This design is optimal for our LiDAR bundle adjustment and scalable mapping, as further elaborated in Sections IV and V.
@@
We also introduce information matrices to describe uncertainties of extracted lines and planes. We define $\sqrt{\mathbf{\Lambda}}=\mathbf{I}\cdot(\sqrt{N/m}/\sigma)$ as the square root of the information matrix, where $(\sqrt{\mathbf{\Lambda}})^{T}\sqrt{\mathbf{\Lambda}}=\mathbf{\Lambda}$ and $m$ is the number of observation points for a landmark. The term $N$ represents the number of raw points in line or plane segments. The variance $\sigma$ varies across different semantic categories. Specifically, $\sigma$ is determined by the flatness of point cloud. We set $\sigma_{r}=0.1,\sigma_{b}=0.2,\sigma_{p}=0.3$ for roads, buildings, and poles in the experimental validation.